OpenStudy (anonymous):

^3 square root of 64y^9

OpenStudy (kc_kennylau):

\[\sqrt[3]{64y^9}=\sqrt[3]{4^3(y^3)^3}=4y^3\]

OpenStudy (anonymous):

thank you